# Augmenting Sequential Recommendation with Balanced Relevance and Diversity

### v2

@@ Introduction
The base mixup can only create samples in linear space. Inspired by the nonlinear mixup (Guo 2020), we attempt to assign different weights to every parameter in $H_{u}$. However, this mixup strategy incurs significant computational overhead. Thus, we further decompose this process into Item-wise mixup and Feature-wise mixup. Given a batch of sequences $\left\{s_{u}\right\}_{u=1}^{N}$, we can obtain the corresponding batch of representations $\left\{H_{u}\right\}_{u=1}^{N}\in R^{B\times N\times D}$, where $B$, $N$ and $D$ are the batch size, maximum sequence length, and embedding dimensions, respectively. We start by shuffling $\left\{H_{u}\right\}_{u=1}^{N}$ from the batch perspective:
$\left\{H_{u}^{\prime}\right\}_{u=1}^{N}=\operatorname{Shuffle}\left\{H_{u}\right\}_{u=1}^{N}.$ (10)
For Item-wise Nonlinear Mixup, we sample a mixup weights matrix $\Lambda_{I}\in R^{N\times L}$ from the $\operatorname{Beta}(\alpha,\alpha)$, where $L$ is the predefined maximum sequence length. The mixup is performed as follows:
$\left\{H_{u}^{Out}\right\}_{u=1}^{N}=\Lambda_{I}\circ\left\{H_{u}\right\}_{u=1}^{N}+(1-\Lambda_{I})\circ\left\{H_{u}^{\prime}\right\}_{u=1}^{N},$ (11)
where $\circ$ denotes the Hadamard product. For Feature-wise Nonlinear Mixup, we sample a mixup weights matrix $\Lambda_{F}\in R^{N\times D}$ from the $\operatorname{Beta}(\alpha,\alpha)$ and perform similar operation:
$\left\{H_{u}^{Out}\right\}_{u=1}^{N}=\Lambda_{F}\circ\left\{H_{u}\right\}_{u=1}^{N}+(1-\Lambda_{F})\circ\left\{H_{u}^{\prime}\right\}_{u=1}^{N}.$ (12)

### v3

@@ Introduction
The base mixup can only create samples in linear space. Inspired by the nonlinear mixup (Guo 2020), we attempt to assign different weights to every parameter in $H_{u}$. However, this mixup strategy incurs significant computational overhead. Thus, we further decompose this process into Item-wise mixup and Feature-wise mixup. Given a batch of sequences $\left\{s_{u}\right\}_{u=1}^{B}$, we can obtain the corresponding batch of representations $\left\{H_{u}\right\}_{u=1}^{B}\in R^{B\times N\times D}$, where $B$, $N$ and $D$ are the batch size, maximum sequence length, and embedding dimensions, respectively. We start by shuffling $\left\{H_{u}\right\}_{u=1}^{B}$ from the batch perspective:
$\left\{H_{u}^{\prime}\right\}_{u=1}^{B}=\operatorname{Shuffle}\left\{H_{u}\right\}_{u=1}^{B}.$ (10)
For Item-wise Nonlinear Mixup, we sample a mixup weights matrix $\Lambda_{I}\in R^{B\times N}$ from the $\operatorname{Beta}(\alpha,\alpha)$, where $N$ is the predefined maximum sequence length. The mixup is performed as follows:
$\left\{H_{u}^{Out}\right\}_{u=1}^{B}=\Lambda_{I}\circ\left\{H_{u}\right\}_{u=1}^{B}+(1-\Lambda_{I})\circ\left\{H_{u}^{\prime}\right\}_{u=1}^{B},$ (11)
where $\circ$ denotes the Hadamard product. For Feature-wise Nonlinear Mixup, we sample a mixup weights matrix $\Lambda_{F}\in R^{B\times D}$ from the $\operatorname{Beta}(\alpha,\alpha)$ and perform similar operation:
$\left\{H_{u}^{Out}\right\}_{u=1}^{B}=\Lambda_{F}\circ\left\{H_{u}\right\}_{u=1}^{B}+(1-\Lambda_{F})\circ\left\{H_{u}^{\prime}\right\}_{u=1}^{B}.$ (12)